OpenStudy (anonymous):

Two normal dice, one blue and one red rolled. the two results added together. What is the probability of getting a sum of 6, given that the blue die is even.

OpenStudy (kropot72):

The sample space has 36 possible combinations of numbers. These can be set out in column form as follows: 6,6 5,6 4,6 3,6 2,6 1,6 6,5 5,5 4,5 3,5 2,5 1,5 6,4 5,4 4,4 3,4 2,4 1,4 6,3 5,3 4,3 3,3 2,3 1,3 6,2 5,2 4,2 3,2 2,2 1,2 6,1 5,1 4,1 3,1 2,1 1,1

OpenStudy (anonymous):

Yes because you have two selections, its 6C1 x 6C1 so you have 36 total combinations.

OpenStudy (kropot72):

However since we are given that the blue die is even, and taking the first of the pairs as the blue die result, there are 3 columns that apply. The sample space is just 18 possible combinations. How many of these sum to 6?

OpenStudy (anonymous):

I just solved it then, P(blue and even) = 36/2 = 18 possible combinations

OpenStudy (anonymous):

then p(sum of 6) = 5/36 however of those 5, two are even and blue. therefore = 2/18

OpenStudy (kropot72):

Of the possible 18 outcomes 2,4 and 4,2 add to 6. The probability is therefore 2/18 = 1/9

OpenStudy (anonymous):

Ok yes that is the right answer. How about P(Total of 5|at least one is even)

OpenStudy (anonymous):

I'm good with algebra, calculus etc is fairly easy for me, but i suck at probability.

OpenStudy (kropot72):

Look at the each of the 36 possible combinations and write down those that total 5 with one either 2 or 4.